The base of a 16ft ladder is 4ft from a wall. How high on the wall does the ladder reach?
Answer by checkley77(12844)

You have a right triangle.
a^2+b^2=c^2
4^2+b^2=256
16+b^2=256
b^2=256-16
b^2=240
b=sqrt240
b=15.49 ft. ans. for the height of the ladder up the wall.